OpenStudy (cmtboy2):

Which equations have a negative solution? Choose all answers that are correct. a.-w/2=11 b.-1/6*=-3 c.-4=y/(-10) -3/4z=8

OpenStudy (cmtboy2):

will fan and madel

OpenStudy (dustinrathke24):

a.-w/2 x 2=11 x 2 -w=22 w=-22

OpenStudy (dustinrathke24):

What's the * for on b?

OpenStudy (cmtboy2):

times

OpenStudy (dustinrathke24):

Times what, you have 6*...

OpenStudy (cmtboy2):

that the queshtin

OpenStudy (dustinrathke24):

There's something missing.

OpenStudy (cmtboy2):

no

OpenStudy (dustinrathke24):

However, the rest are as follows: c. -4=y/-10 y/-10 x -10 = -4 x -10 y=40 So that's positive.

OpenStudy (dustinrathke24):

a.-w/2 x 2=11 x 2 -w=22 w=-22 So that's negative.

OpenStudy (dustinrathke24):

Ohhhh, did you mean -1/6x=-3?

OpenStudy (dustinrathke24):

b.-1/6x=-3 -1*1/6x=-3 -1*1/6x*6x=-3*6x -1=-18x x=1/18 So that's positive.

OpenStudy (dustinrathke24):

d. -3/4z=8 -3*1/4z=8 -3*1/4z*4z=8*4z -3=32z z=-3/32 So that's negative.

OpenStudy (dustinrathke24):

The x on -1/6x is a variable, not a multiplication sign.